Two stationary charged objects are 20.0 cm apart from each other. Object B is (-2.6mC) and object L is (-3.4mC). What is the electric field (provide the magnitude and direction) they generate at the midpoint between them?
1
Expert's answer
2019-04-15T09:47:33-0400
"E_B=\frac{kq_B}{r^2}=\frac{(9 \cdot 10^9)(2.6 \cdot 10^{-3})}{0.1^2}=2.34 \cdot 10^9 \frac{N}{C}"

Directed to the object B.


"E_L=\frac{kq_L}{r^2}=\frac{(9 \cdot 10^9)(3.4 \cdot 10^{-3})}{0.1^2}=3.06 \cdot 10^9 \frac{N}{C}"

Directed to the object L.

Net electric field:


"E=E_L-E_B=3.06 \cdot 10^9-2.34 \cdot 10^9=7.2 \cdot 10^8 \frac{N}{C}"

Directed to the object L.